Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

The probability of winning game A is X and the probability

Expert replies
by BTGmoderatorDC » Mon Oct 08, 2018 6:02 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

The probability of winning game A is X and the probability of winning game B is Y. When playing a single round of each of the two games, what is the probability of winning exactly once?

A. X-YX
B. XY
C. X+Y
D. X+Y-2XY
E. X^2Y^2-XY^2-X^2Y

OA D

Source: Economist GMAT
Join the discussion
Source: — Problem Solving |

by Jay@ManhattanReview » Tue Oct 09, 2018 4:29 am
BTGmoderatorDC wrote:The probability of winning game A is X and the probability of winning game B is Y. When playing a single round of each of the two games, what is the probability of winning exactly once?

A. X-YX
B. XY
C. X+Y
D. X+Y-2XY
E. X^2Y^2-XY^2-X^2Y

OA D

Source: Economist GMAT
Given:

1. The probability of winning game A is X, and
2. The probability of winning game B is Y

Thus,

The probability of NOT winning game A = 1 – X
The probability of NOT winning game B = 1 – Y

Thus, the probability of winning exactly once = the probability of winning either A or B but not both the games

= [(Probability of winning game A)*(Probability of NOT winning game B)] + [(Probability of winning game B)*(Probability of NOT winning game A)]

= X*(1 – Y) + Y*(1 – X)
= X – XY + Y – XY
= X + Y – 2XY

The correct answer: D

Hope this helps!

-Jay
_________________
Manhattan Review GMAT Prep

Locations: New York | Bangkok | Abu Dhabi | Rome | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion

by Scott@TargetTestPrep » Fri Oct 12, 2018 6:55 am
BTGmoderatorDC wrote:The probability of winning game A is X and the probability of winning game B is Y. When playing a single round of each of the two games, what is the probability of winning exactly once?

A. X-YX
B. XY
C. X+Y
D. X+Y-2XY
E. X^2Y^2-XY^2-X^2Y
P(winning exactly once) = 1 - P(winning zero times) - P(winning both times)

P(wWinning exactly once) = 1 - (1 - X)(1 - Y) - XY = 1 - (1 - X - Y + XY) - XY = X + Y - 2XY

Alternate Solution:

There are two ways of winning exactly once, either (Win A, Lose B) OR (Lose A, Win B). The probability of winning A is X, so the probability of losing A is (1 - X). Similarly, the probability of winning B is Y, and so the probability of losing B is (1 - Y).

P(Win A, Lose B) = X*(1 - Y) = X - XY

P(Lose A, Win B) = (1 - X)*Y = Y - XY

The probability that either of these events happens is the sum of their individual probabilities, so P(Winning exactly once) = X - XY + Y - XY = X + Y - 2XY

Answer: D

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

BTGmoderatorDC wrote:The probability of winning game A is X and the probability of winning game B is Y. When playing a single round of each of the two games, what is the probability of winning exactly once?

A. X-YX
B. XY
C. X+Y
D. X+Y-2XY
E. X^2Y^2-XY^2-X^2Y

Source: Economist GMAT
$$? = P\left( {{\text{win}}\,\,{\text{just}}\,\,{\text{A}}\,\,{\text{or}}\,\,{\text{win}}\,\,{\text{just}}\,\,{\text{B}}} \right)\,\,\,\, = \,\,\,\,P\left( {{\text{win}}\,\,{\text{A}}\,{\text{or}}\,\,{\text{B}}} \right) - P\left( {{\text{win}}\,\,{\text{both}}} \right)$$
$$P\left( {{\rm{win}}\,\,{\rm{both}}} \right) = xy\,\,\,\,\,\,\,\,\left[ {\,x,y\,\,\,{\rm{constants}}\,\,\,\, \Rightarrow \,\,\,\,{\rm{independency}}\,} \right]$$
$$P\left( {{\text{win}}\,\,{\text{A}}\,{\text{or}}\,\,{\text{B}}} \right) = x + y - xy\,\,\,\,\,\left[ {\,{\text{simplifier}}\,} \right]$$
$$? = x + y - 2xy$$

Image

This solution follows the notations and rationale taught in the GMATH method.

Regards,
Fabio.


"Kolmogorov – Poincaré – Gauss – Euler – Newton, are only five lives separating us from the source of our science".
(Vladimir Arnold)
Last edited by fskilnik@GMATH on Sat Oct 13, 2018 1:30 pm, edited 2 times in total.
Fabio Skilnik :: GMATH method creator ( Math for the GMAT)
English-speakers :: https://www.gmath.net
Portuguese-speakers :: https://www.gmath.com.br
Join the discussion

by [email protected] » Fri Oct 12, 2018 4:18 pm
Hi All,

We're told that the probability of winning Game A is X and the probability of winning Game B is Y. We're asked when playing a SINGLE round of EACH of the two games, what is the probability of winning EXACTLY ONCE. This question can be solved in a couple of different ways, including by TESTing VALUES.

IF... X = .4 AND Y = .5
then there is a 40% chance to win Game A and a 60% chance to lose Game A
and a 50% chance to win Game B and a 50% chance to lose Game B.

The probability of winning A and losing B = (.4)(.5) = .2
The probability of winning B and losing A = (.5)(.6) = .3
Total probability of winning just one Game = .2 + .3 = .5

Thus, we're looking for an answer that equals .5 when X = .4 and Y = .5 There's only one answer that matches....

Final Answer: D

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion